OpenStudy (anonymous):

Diana received 65 points on a project for school. She can make changes and receive three-tenths of the missing points back. She can do this 10 times . Create the formula for the sum of this geometric series, and explain your steps in solving for the maximum grade Diana can receive.

OpenStudy (anonymous):

@NathalyN

OpenStudy (nathalyn):

r = 3/10

OpenStudy (anonymous):

Yes

OpenStudy (nathalyn):

okay that is your answer!?

OpenStudy (anonymous):

How?

OpenStudy (nathalyn):

10 is the naswer

OpenStudy (anonymous):

But it wants you to make a formula. @n

OpenStudy (nathalyn):

100 - 65 = 35 (assuming this is 100) she lost 35 points, she can get 3/10 of those points back '35' is the missing point * 310 = 35∗31∗10=10510=10.5 65 + 10.5 = 75.5

OpenStudy (nathalyn):

does that make more sens4e3?????

OpenStudy (anonymous):

It wants a formula though

OpenStudy (anonymous):

@NathalyN

OpenStudy (is3535):

I'm going to try and make sense in this, I only have a few minutes left. 100 - 65 = 35 (assuming this is 100) she lost 35 points, she can get 3/10 of those points back '35' is the missing point * \(\ \dfrac{3}{10} \) = \(\ \dfrac{35 * 3}{1 * 10} = \dfrac{105}{10} = 10.5 \) 65 + 10.5 = 75.5

OpenStudy (is3535):

yes, the first time she missed something, and she got a 65/100, she is resubmitting to gain back 3/10ths of the missing points, 35

OpenStudy (anonymous):

So 310 = 35∗31∗10 =10510 =10.5 65 + 10.5 = 75.5 is the answer @is3535

OpenStudy (is3535):

yes 75.5 is your answer

OpenStudy (nathalyn):

that is what i said!????

OpenStudy (is3535):

not late can i have a medal

OpenStudy (anonymous):

Okay thanks

OpenStudy (is3535):

can i have a medal

OpenStudy (anonymous):

@NathalyN it written weird and I think thats why I didnt get it because I couldn't see where the fractions were. I gave you a medal and fan on the last question